1. The alloy aluminum magnesium boride (called BAM) has one of the lowest coefficients of friction: 0.02. A ramp with a 30.0° incline has a BAM coating. A box with a mass of 122 kg slides down the ramp.

a) Calculate the acceleration of the box

b) How does this value compare to the acceleration of the box if the ramp were truly frictionless?



1
Expert's answer
2021-05-12T08:37:44-0400

a)


"a=(\sin{30}-\mu \cos{30})g\\a=(\sin{30}-0.02\cos{30})9.8=4.73\frac{m}{s^2}"

b)


"a=(\sin{30})g\\a=(\sin{30})9.8=4.9\frac{m}{s^2}"